OpenStudy (refrac532):

Help please, my physics teacher neglected to teach me hang time and I can't find any good results online. Calculate the hang time of an athlete who jumps a vertical distance of 0.75 meters. *Note that this is 8th grade introductory physics

OpenStudy (astrophysics):

Hang time is just the time he was in the air so we can use the kinematics formula \[d = v_i+1/2at^2\] where a = acceleration due to gravity which is 9.81 m/s^2

OpenStudy (astrophysics):

the initial velocity is 0, and d = 0.75, solve for t

OpenStudy (astrophysics):

@refrac532

OpenStudy (refrac532):

So, rearrange the equation to solve for time?

OpenStudy (astrophysics):

Correct

OpenStudy (astrophysics):

Solve for \[d = 1/2at^2\] since initial velocity is 0

OpenStudy (astrophysics):

I should've used g instead of a, but it doesn't matter just note you're using gravity

OpenStudy (refrac532):

Um, so my teacher says that 10 m/s^2 is the gravitational acceleration for Earth (yes, I know it's 9.8 m/s^2), so would the equation be\[0.75 = 5m/s^2t^2\]?

OpenStudy (refrac532):

I'm sorry @Astrophysics , I'm still a little confused about how to isolate t

OpenStudy (astrophysics):

I like to do the algebra first then plug in the numbers, but it's cool if you plug in the numbers in this case it made it easier! So we have \[\frac{ (0.75m) }{ 5 m/s^2 } = t^2\] we divide both sides by 5 m/s^2 to cancel out the 5 m/s on the right, any idea how to get rid of the square on the t? Hint: \[\sqrt{x^2} = x\]

OpenStudy (refrac532):

So it would be \[\sqrt{\frac{ 0.75m }{ 5m/s^2 }} = t\] right?

OpenStudy (astrophysics):

Exactly!

OpenStudy (refrac532):

I got about 0.387 seconds. @Astrophysics

OpenStudy (astrophysics):

Since the question is in 2 sig digs, lets stick with 2! So 0.39 s :)

OpenStudy (refrac532):

Okay, thanks a lot!

OpenStudy (astrophysics):

Np